Prev Next

SW&IT Meeting Minutes 2018-09-17

Attendees: Andrew, David, Igor, Kellee, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

News from management

  • xxx

Previous Action Items

  • Vsphere (DE)
  • power plan update draft (DE)
  • Luci H/O doc (IS)
  • Luci telemetry
    • web-io cold heads state for vibration (IS)
    • luci telemetry in git

Transition plan

Handover steps

  1. primary + secondary do self learning, gather information, read wiki
  2. primary organize teamviewer/zoom walk through with Kellee and secondary
  3. primary draft techtalk w. input from secondary
  4. techtalk review by Kellee
  5. primary does a team presentation

Please update this table to indicate progress.


Application(s) Primary Secondary Week
OAC
AC (2)
Petr (2)
1
AzCam
AC (2)
Petr (2)
2
DIMM
AC (2)
Petr (2)
2
Weatherstation
AC (2)
Petr 3
FACSUM,WEATHERSTATION
AC (2)
Matthieu 4
ALH Igor(1/2)
Matthieu (2)
1
INDI Igor Matthieu 3
GCS
Petr (3)
Matthieu (2)
1
Telemetry
Petr (3)
Matthieu (2)
2
OVMS
Petr (3)
Matthieu (2)
3
Status Server Stephen (1)
Matthieu (1)
2
Telemetry cleanup Stephen Petr 3
LBC
Yang (1/2)
Igor(1/2) 2
MODS
Yang (1)
Igor 2
Telem-VIZ Yang Stephen 3
particle monitors Yang Petr 4

Open topics

Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.

This week in github

Team reports

PK

  • cleaned office white table - can only highly recommend
  • GCS - WFS related work, resulting in GCS (re) desing ideas (IT 7388, 7407)
  • discovered "git show" command
  • TCS 2018B branch and setup
  • vim-airline plugin
  • (trivial) Dockerfile to build and run TCS in Docker

next:

  • GCS etc. HO presentations
  • GCS future ideas
  • git grep command
  • mountain night trip?

AC

Last week:
  • Prepared OAC and AzCam tech talks
  • DIMM knowledge transfer with KS
  • Troubleshot ARGOS deployment on sx-lgsw (dependency issues like cfitsio, more...)

This week:
  • Investigate MPIA repo inconsistencies
  • OAC, AzCam, and DIMM tech talks
  • ARGOS deployment
  • IT 7388 code review

KS

Last week
  • LBC Windows environment documentation on the wiki, including screenshots.
  • IT 7380 discussions - LBC is zeroing the proper motion values in the TCS interface.
  • LBC discussion with Yang, Igor, and Olga - walked through consolidated log, describe the error handling between subsystems, discussions about power control.
  • Updated ALH and Status Server with both MODS after installation on the telescope.
  • Weatherstation discussion with Andrew.

This week
  • Vacation Monday/Tuesday.
  • Continue going through all my notes to pass on whatever's important.
  • Review handover presentations?

YZ

  • Working on transition to Kellee's LBC responsibility.
    • Met with Kellee, Igor, and Olga another time.
    • Going through wiki and editing notes.
  • Getting started with MODS (but not much yet).
  • Telescope collimation seemed to be okay, but still needs work on current collimation tools.
  • Continue working on documenting pointing and collimation procedures.
  • Continue working on Python package and modules.
  • Need to start gather more info on Al's AO project.

MB

  • management OC, WPC
  • Handover activities: StatusServer
  • Labor Reporting stats
  • Shark-NIR schema review
  • TMT ESW review
  • LBTI/single mode fiber follow up
  • jenkins/docker review with SH
  • bug fix IMS

next:
  • obs workstation upgrade
  • power upgrade plan
  • IT to github issues
  • TMT ESW review
  • jumbo / Soul

DE

Last week:
  • VMWare upgrade work
    • Have upgrade packages
    • Will require downtime
  • UPS vendor chat
  • Peter monitors arrived in good condition
This week:
  • Peter computer setup
  • Chat about server room power (Wednesday?)

IS

Last Week:
  • HandOver
    • LBC:
      • Install of CLion
      • went through some of the code as part of #7380
      • meeting with Kellee and Yang
    • LUCI
      • Send handover document to Matthieu
      • Received corrections from Matthieu
  • PIT
    • Fix small bug in PDF table and sent final release
  • LUCI RTD work
  • OT/LBC #7604
    • Fixed development webserver. Thank you David Edmeades for the help!!
    • Bug affects calibrations in archive, calibration files are available to all partners and not just to the partner that took the observation. Script keywords need to be "CALIBRATION" instead of partner for it to be widely available.
    • waiting for Jenny to confirm testing.
This Week:
  • LUCI development of AFC RTD, almost done need to sync up with Dave for testing
  • Handover
  • Complete LUCI handover document updates

SH

Last week:

  • Thursday trip up the mountain was mostly useless

This week:

  • Meeting to discuss the LBTI fiber runs -- finding a time is difficult.
  • Look at statserv more in depth.
  • Start playing with Docker cluster again.
  • Talk to MGIO about networking, camera at basecamp.
  • Fix up LBTI networking after talking to Paul (change VLAN and IP of hosts).

Topic revision: r22 - 10 Dec 2018, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback